### v1

# AED: Automatic Discovery of Effective and Diverse Vulnerabilities for Autonomous Driving Policy with Large Language Models

Abstract
Assessing the safety of autonomous driving policy is of great importance, and reinforcement learning (RL) has emerged as a powerful method for discovering critical vulnerabilities in driving policies. However, existing RL-based approaches often struggle to identify vulnerabilities that are both effective—meaning the autonomous vehicle is genuinely responsible for the accidents—and diverse—meaning they span various failure types. To address these challenges, we propose AED, a framework that uses large language models (LLMs) to Automatically discover Effective and Diverse vulnerabilities in autonomous driving policies. We first utilize an LLM to automatically design reward functions for RL training. Then we let the LLM consider a diverse set of accident types and train adversarial policies for different accident types in parallel. Finally, we use preference-based learning to filter ineffective accidents and enhance the effectiveness of each vulnerability. Experiments across multiple simulated traffic scenarios and tested policies show that AED uncovers a broader range of vulnerabilities and achieves higher attack success rates compared with expert-designed rewards, thereby reducing the need for manual reward engineering and improving the diversity and effectiveness of vulnerability discovery.
I INTRODUCTION
Autonomous driving is a safety-critical domain where even minor decision-making flaws can lead to severe accidents [1]. One approach to assessing the safety of autonomous driving policies is vulnerability discovery, which systematically uncovers scenarios where autonomous driving policy fails or behaves undesirably [2, 3]. By controlling the behaviors of other traffic participants, vulnerability discovery create adversarial driving scenarios where the tested driving policy is exposed to hazardous conditions. Due to the rarity of such failure cases in real-world road tests, automated vulnerability discovery methods have been developed in simulation environments, where reinforcement learning (RL) emerges as an efficient approach that trains adversarial agents to induce accidents of the tested policy [4, 5, 6, 7].
Figure 1: Illustration of effective and diverse vulnerabilities. An effective vulnerability is an accident that is caused by the false decision-making of the tested policy. Diverse vulnerabilities refer to different accident types that are consistent with human understanding and regulations.
However, there are two key challenges in leveraging RL for vulnerability discovery. First, it is hard to automatically identify effective vulnerabilities that require determination of accident responsibility. When an accident occurs, it is essential to attribute the failure to the autonomous vehicle’s (AV) decision-making or those of other traffic participants. An effective vulnerability is one where the tested policy is responsible for the accident, revealing a genuine deficiency in its driving policy [8, 9, 7]. Accident responsibility, however, is complex to assess, often requiring human experts with domain knowledge in real-world settings [10, 11]. Similarly, in simulation-based RL training, designing a suitable reward function for responsibility attribution also demands substantial domain expertise [12, 13, 14], posing a challenge to the automated discovery of effective vulnerabilities.
The second challenge in RL-based vulnerability discovery is ensuring diverse failure cases. Since a safety-critical flaw can arise from various accident types, it is crucial to comprehensively explore the space of possible failures [9, 15, 7]. Traditional RL methods, however, tend to converge on the most probable failure case, leading to a lack of diversity in discovered vulnerabilities [16]. Existing approaches to diversity-aware RL generally fall into two types. The first type is optimizing a general diversity metric alongside task objectives. However, these methods often lead to vulnerabilities with uninterpretable differences that are not consistent with human understanding and regulations. The second type is incorporating heuristic-driven diversity measures tailored to specific tasks, which rely heavily on expert knowledge and struggle to generalize across different scenarios. These limitations hinder the interpretability and automation of diverse vulnerability discovery, leaving potential safety risks unsolved.
To address the aforementioned challenges, we propose AED, a framework that leverages large language models (LLMs) to Automatically discover Effective and Diverse vulnerabilities in autonomous driving policies. Our framework consists of three key steps. First, an LLM is prompted with the environment and task descriptions to automatically design reward functions for vulnerability discovery. Then, to enhance the diversity of vulnerabilities, we use the LLM to consider a broad range of accident scenarios and train adversarial policies for different types in parallel. Finally, to further improve the effectiveness of the discovered vulnerabilities, we use preference-based reinforcement learning to train a reward model that filters ineffective accidents. By combining automatic reward design, diverse accident generation, and preference-based effectiveness enhancement, our framework automatically identifies a broad spectrum of effective and diverse vulnerabilities, improving the comprehensiveness of AV safety evaluation.
To evaluate the performance of our framework, we apply our method to discover vulnerabilities of both planning-based policies and learning-based policies across different traffic scenarios, involving different numbers of adversarial vehicle agents. Experiment results show that our method can successfully identify and distinguish a diverse set of different vulnerabilities. Moreover, our method is able to achieve higher attack success rates with automatically generated rewards than policies with expert-designed rewards. These results show that our proposed framework can discover a diverse set of effective vulnerabilities in various tested policies, significantly reducing the reliance on manually designed reward functions and diversity metrics.
In summary, the contributions of this paper are: (1) We propose to use LLMs to automatically discover a diverse set of potential vulnerabilities, enhancing the overall safety of a driving policy. (2) We propose a preference-based learning method to filter ineffective accidents and encourage the identification of effective vulnerabilities. (3) We demonstrate our approach through extensive experiments and analysis across various traffic environments and tested driving policies, demonstrating its robustness and generalizability.
II RELATED WORK
Figure 2: Overview of our proposed AED framework that uses large language models (LLMs) to automatically discover effective and diverse vulnerabilities in autonomous driving policies. Our framework first utilizes an LLM to automatically design reward for vulnerability discovery. Then we use the LLM to consider different accident types and generate a diverse set of accidents in parallel. Finally, we use preference-based learning to enhance the effectiveness of each accident type. Combining the three components leads to our framework for automatic, effective, and diverse vulnerability discovery.
II-A Diverse Vulnerability Discovery
Adversarial RL has emerged as a popular approach to purposely generating safety-critical scenarios to efficiently find vulnerabilities where autonomous driving policies make mistakes. Kuutti et al. [4] use Advantage Actor Critic (A2C) [17] to induce rear-end collision. Some works [18, 5, 19] use Deep Q-Network [20] to generate corner cases to elicit AV mistakes. Chen et al. [6] use Deep Deterministic Policy Gradient (DDPG) [21] to create challenging lane-change conditions. Adaptive Stress Testing [9] incorporates a Responsibility-Sensitive Safety (RSS) [10] reward into Trust Region Policy Optimization (TRPO) [8] to classify AV failures with a single adversarial agent. Some works [22, 7] utilize Multi-agent DDPG [23] and Multi-agent Proximal Policy Optimization (PPO) [24] respectively to coordinate multiple adversarial agents. Some other studies [25, 26] use genetic algorithms to evolve adversarial agents to attack AV. However, these methods rely on manually designed reward functions, requiring extensive domain expertise to balance responsibility attribution and behavioral realism, which limits generalization across diverse traffic environments.
To improve the diversity of discovered vulnerabilities, existing methods can be categorized into general and heuristic methods. General methods maximize information-theoretic objectives to encourage trajectory diversity [16] or measure geographic and semantic similarity between failure trajectories [27, 9]. However, these methods can lack interpretability for real-world accidents. Heuristic methods incorporate expert knowledge or discovered datasets to enhance diversity. Chen et al. [6] cluster multiple scenarios using DP-Means [28], while VDARS [7] proposes intrinsic rewards to promote novel failures. However, these methods still rely on manual design and domain-specific assumptions. Our work overcomes these limitations by using LLMs to automatically generate diverse reward functions for different vulnerability types, improving adaptability and reducing reliance on handcrafted heuristics.
II-B Reward Design with LLMs
Reward design is a fundamental challenge in reinforcement learning (RL), as it requires crafting reward functions that both reflect the intended behaviors and offer effective learning signals. Traditional reward engineering is often labor-intensive, relying heavily on domain expertise and iterative trial-and-error processes [29, 30]. With the rise of large language models (LLMs) pretrained on extensive human knowledge, their capabilities in reasoning, planning, and code generation have opened new possibilities for addressing core RL problems [31, 32, 33].
Among recent advances, utilizing LLMs for automatic reward modeling has become a promising direction. Some studies employ LLMs to interpret the environment and agent behavior, providing intrinsic rewards based on scenario understanding [12, 34, 35, 36]. Others focus on leveraging LLMs to generate executable reward code, often incorporating self-refinement mechanisms to iteratively improve the reward functions [13, 14, 37, 38]. Notably, Eureka [13] introduces an automatic reward design framework based on self-reflection. This approach leverages LLMs to generate executable reward functions using the environment’s source code and iteratively refines these functions based on evaluation results obtained from human-designed sparse rewards. Building on this concept, In-Context Preference Learning (ICPL) [14] extends the framework to tasks where human-designed rewards are unavailable, enabling LLMs to directly learn reward functions from human preferences. Our work extends ICPL to multi-agent and dynamic traffic scenarios, where reward design should account for complex agent interactions and responsibility attribution.
II-C Preference-Based Reinforcement Learning
Preference-based reinforcement learning has been proposed to learn directly from an expert’s preference instead of handcrafted numeric reward functions, thereby reducing the reliance on extensive domain-specific prior knowledge. Studies [39, 40, 41, 42, 43, 44, 45] have explored RL from human feedback (RLHF) in the form of trajectory preference queries or score ranking to train a reward model. However, continuously querying human labels during online training is often impractical due to the associated cost and latency. To address this limitation, some studies have proposed RL from artificial intelligence feedback (RLAIF) [46]. Some works [47, 48] use Vision-Language-Models (VLMs) to align vision-based trajectory inputs with task-specific language descriptions. Some works [49, 50, 51] generate intrinsic reward functions using preference feedback from LLMs based on text descriptions of states or trajectories. Our work generates preference data pairs automatically using datasets produced by parallel adversarial policies. This enables scalable preference-based learning without numerous labelling while capturing diverse vulnerabilities under multi-agent interactions.
III PRELIMINARY
III-A Partially Observed Markov Decision Process
Our MARL framework can be described as a Partially Observed Markov Decision Process (POMDP). defined by a tuple $\mathcal{M}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{N}=\{1,2,\cdots,N\}$ is the set of agents, $\mathcal{S}$ is the state space encompassing all possible states of the N agents, $\mathcal{A}$ is the shared action space, $\mathcal{O}$ denotes each agent’s observation space, $\mathcal{P}$ is the transition probability function, $\mathcal{R}$ is the joint reward function with $\gamma$ is the reward discount factor. At each step, agent $i$ receives an observation $o_{i}=\mathcal{O}(s_{i})$ where $s_{i}\in\mathcal{S}$ is the local state. The agent then takes an action $a_{i}$ according to the shared policy $\pi_{\theta}(a_{i}|o_{i})$, where $\theta$ represents the shared parameters. In the next step, the state progresses to $s_{i}^{\prime}$ according to the transition probability $p(s_{i}^{\prime}|s_{i},a_{i})\in\mathcal{P}$ and the agent $i$ receives a reward $r_{i}=\mathcal{R}(s_{i}|a_{i})$. For homogeneous agents, each agent $i$ optimizes the discounted cumulative rewards $J(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t}\gamma^{t}r_{t}\right]$, where a trajectory $\tau$ denotes a state-action-reward triplets sequence, i.e., $\tau=\{(s_{t},a_{t},r_{t})\}_{t}$.
III-B Effective and Diverse Vulnerability Discovery
Our objective is to discover effective and diverse vulnerabilities in a fixed, pre-trained autonomous driving policy by training adversarial policies within the previously defined POMDP framework. While the tested driving policy remains unchanged, adversarial vehicle agents are trained to induce failure scenarios that expose weaknesses in the tested policy. Specifically, the adversarial policy is optimized to maximize the rates of effective vulnerability—defined as those for which responsibility can be attributed to the tested policy. The diversity is evaluated by the number of distinct vulnerability types, and the effectiveness is evaluated by the proportion of vulnerabilities discovered vulnerabilities attributable to the tested driving policy’s wrong decisions.
IV METHODOLOGY
To automatically find effective and diverse vulnerabilities, we propose the AED framework that consists of three components including automatic reward design, diverse accident generation, and preference-based effectiveness enhancement. The overview of our framework is shown in Fig. 2. The first automatic reward design component prompts an LLM with the environment and task descriptions to automatically generate reward function samples and select the best one. Then, to enhance the diversity of vulnerabilities, we propose a diverse accident generation component that prompts the LLM to consider a broad range of accident types and trains adversarial policies for different types in parallel. Finally, to enhance the effectiveness of each vulnerability, we use preference-based learning to train a reward model that filters ineffective accidents and other accident types. Combining the three components lead to our framework that automatically discovers effective and diverse vulnerabilities.
IV-A Automatic Reward Design
To enable automatic reward design for vulnerability discovery, we extend the In-Context Preference Learning (ICPL) framework [14] to leverage the code-synthesis and in-context learning capabilities of LLMs. In our adaptation, the LLM generates task-specific reward functions based on environment descriptions and human preferences, guiding the discovery of effective accidents in a self-improving manner.
The LLM first receives an environment description and a detailed task specification, then generates $K$ executable reward functions. The environment text outlines the scenario and relevant code-level details, such as the agent’s observation/action space and key parameters or functions useful for reward construction. The specific details of the task description are provided in Sec. IV-B. Each reward function is used to train an agent, resulting in $K$ agents, which are evaluated using human-written criteria for identifying effective accidents. To ensure comprehensive coverage, we also generate accident videos for human evaluation. Based on both metrics and visual inspection, evaluators select the best and worst reward functions. The selected best function typically yields high accident effectiveness across scenarios and serves as a basis for further refinement.
IV-B Diverse Accident Generation
To enhance diversity in vulnerability discovery, the adversarial policy should capture distinct characteristics of various weaknesses in the tested policy. To achieve this, we introduce a Diverse Accident Generation component, which prompts LLMs to generate a set of reward functions that promote vulnerability diversity. By considering a wide range of accident types, LLMs automatically design multiple reward functions in parallel, each tailored to induce specific failure cases in the tested policy.
The process begins with the LLM generating a task description for a given accident type based on the environment specifications. This description includes a high-level summary of the targeted accident scenario, concrete examples, justifications for attributing responsibility to the tested agent, and well-defined criteria for categorizing scenarios within this accident type. Using the component outlined in Sec. IV-A, the LLM then autonomously generates reward functions corresponding to different accident types in parallel.
Once the reward functions are generated, they are integrated into the training of adversarial policies, each optimized for a specific vulnerability of the tested policy. These adversarial policies are then used to collect trajectory data, creating a structured dataset covering diverse failure cases. This dataset serves as a foundation for further analysis and improvement of the tested policy, ensuring a comprehensive evaluation of its robustness across various accident scenarios.
IV-C Preference-Based Effectiveness Enhancement
Reward functions derived from Sec. IV-B for a specific accident type are often imperfect, sometimes capturing unintended accident types or ineffective scenarios. For instance, the reward function for Left Lane Change (Accident Type 1) may capture Right Lane Change (Accident Type 2) and Adversarial Vehicle cut in mistakenly (Accident Type 3). To improve the effectiveness and distinguish diverse accident types, we propose a preference-based effectiveness enhancement component to refine the LLMs-driven reward function. The key idea is to learn a new reward function $\hat{r}$ that prefers trajectories corresponding to the intended accident types over other types. Specifically, $\hat{r}$ is trained to assign higher cumulative rewards to Type 1 than to those from Type 2 or Type 3.
To construct training data for preference learning, we leverage trajectory data generated in parallel for different accident types using their respective LLMs-driven reward functions. We consider the Effective and Type 1 dataset as Positive Data $\sigma^{1}$, and use a combination of the Ineffective or Other Types (Type 2 and Type 3) as Negative Data $\sigma^{2}$.
The probability of preferring a trajectory $\sigma^{i}$ as the intended accident type is assumed to depend exponentially on the accumulated latent reward over the trajectory’s state-action sequence $(o_{t}^{i},a_{t}^{i})$ :
$\hat{P}[\sigma^{1}\succ\sigma^{2}]=\frac{\exp\sum_{t}\hat{r}(o_{t}^{1},a_{t}^{1})}{\exp\sum_{t}\hat{r}(o_{t}^{1},a_{t}^{1})+\exp\sum_{t}\hat{r}(o_{t}^{2},a_{t}^{2})}.$ (1)
Figure 3: Number of vulnerability types discovered across different environments, tested policies, and numbers of adversarial vehicle agents. Each line represents the cumulative number over three training iterations, where the first discovered type is counted as 1, and subsequent new types are weighted by their relevant frequency among effective failures. AED consistently discovers more diverse vulnerabilities than VDARS.
To optimize the reward function $\hat{r}$, we minimize the cross-entropy loss between these predictions and the intended preference. The Positive Data and Negative Data are stored in a database $\mathcal{D}$ of triples $(\sigma^{1},\sigma^{2},\mu)$, where $\mu$ is a judgment label (0 or 1) to indicate whether $\sigma^{1}$ is the preferred accident type. The loss function is defined as:
$\text{loss}(\hat{r})=-\sum_{(\sigma^{1},\sigma^{2},\mu)\in\mathcal{D}}$ $\mu\log\hat{P}[\sigma^{1}\succ\sigma^{2}]+$
$(1-\mu)\log(1-\hat{P}[\sigma^{1}\succ\sigma^{2}]).$ (2)
Since $\mu$ assigns full probability mass to the Positive Data, which means $\mu=1$ across the dataset, the loss simplifies to:
$\text{loss}(\hat{r})=-\sum_{(\sigma^{1},\sigma^{2},\mu)\in\mathcal{D}}\log\hat{P}[\sigma^{1}\succ\sigma^{2}].$ (3)
In this way, the reward function $\hat{r}$ learns to favor Type 1 accident trajectories while effectively distinguishing them from other types. We then incorporate $\hat{r}$ and the original LLMs-driven reward function into RL training. If $\hat{r}$ identifies trajectories as Type 2, the training will be pernalized. This guides RL training to prioritize Accident Type 1 and ultimately improve the attack success rate for this category.
V EXPERIMENT
We evaluate AED in two driving environments—Highway and Roundabout—using the Highway simulator [52]. We consider three types of tested driving policy: two Planning-based methods, Value Iteration (VI) [53] and Robust Value Iteration (RVI) [54], and one RL-based method, Dueling Double Q-Network (D3QN) [55]. To account for the complexity of traffic interactions, we conduct experiments with two and three adversarial vehicle agents in each environment.
V-A Implementation Details
We consider three distinct types of vulnerabilities in our experiments: Left Lane Change (Left), Right Lane Change (Right), and Rear End (Rear) collision scenarios. These categories correspond to the different failure-inducing situations: an accident caused during a left lane change maneuver, a right lane change maneuver, or a rear-end collision. To ensure robustness, each adversarial training is repeated with three different random seeds. Each adversarial policy is trained using the MAPPO algorithm [24] for 10 million steps. We utilize GPT-4o for diverse accident generation.
We compare the performance of our method with the following baselines:
  • VDARS [7]: A multi-agent RL framework that employs handcrafted reward functions to learn the adversarial policy for vulnerability discovery.
  • Parallel Eureka: An ablated version of our method without the preference-based effectiveness enhancement module. This can also be regarded as a parallel version of Eureka for different vulnerability types.
V-B Evaluation of Diversity
To evaluate the capability of AED in discovering diverse vulnerability scenarios, we measure the number of distinct vulnerability types identified across multiple training iterations. For each traffic environment setting—defined by the traffic environment (Highway and Roundabout), the tested policy (VI, RVI, and D3QN), and the number of adversarial vehicle agents (2 and 3)—we conduct three independent training iterations with different random seeds. In each iteration, we collect all effective failure cases. In the first iteration, we record the most frequently occurring effective vulnerability type and assign it a count of 1. In the second and third iterations, we exclude types previously discovered and select the most frequently occurring type among the remaining types. Instead of simply incrementing the count by one, we add the relative frequency of the newly discovered type—its proportion among all effective failures in that iteration—to the cumulative count. In this way, each iteration contributes at most one type, and the final number reflects both the counts and the prominence of distinct vulnerability types discovered. For VDARS, each adversarial policy is trained using the same handcrafter reward function across all iterations. For AED, the adversarial policy in each iteration is trained using one of the three different reward functions, which are automatically generated by LLMs in parallel and targeted at different failure types (i.e., Left, Right, and Rear).
As shown in Fig. 3, AED consistently discovers a greater number of distinct vulnerability types compared to VDARS across all tested scenarios. In both the Highway and Roundabout environments, AED reaches or closely approaches the maximum of three types within three training iterations, regardless of the tested policy and the number of adversarial vehicle agents. In contrast, VDARS typically discovers only one or two types, even after multiple iterations. This discrepancy highlights the limited exploration capability of using a single handcrafted reward function in VDARS. On the other hand, AED benefits from multiple LLMs-driven reward functions that promote the exploration of different failure scenarios, enabling the adversarial agents to expose a broader range of policy weaknesses.
V-C Evaluation of Effectiveness
2 Adversaries 3 Adversaries
VI RVI D3QN VI RVI D3QN
Highway Left VDARS $34.24\%_{(0.09)}$ $61.88\%_{(0.52)}$ $25.81\%_{(0.07)}$ $44.38\%_{(0.14)}$ $79.51\%_{(0.10)}$ $58.24\%_{(0.18)}$
Eureka $97.22\%_{(0.08)}$ $100\%_{(0.00)}$ $73.89\%_{(0.55)}$ $97.22\%_{(0.08)}$ $97.78\%_{(0.08)}$ $32.22\%_{(0.57)}$
AED $\mathbf{100\%}_{(0.00)}$ $\mathbf{100\%}_{(0.00)}$ $\mathbf{90.00\%}_{(0.04)}$ $\mathbf{100\%}_{(0.00)}$ $\mathbf{100\%}_{(0.00)}$ $\mathbf{88.89\%}_{(0.05)}$
Right VDARS $39.56\%_{(0.11)}$ $11.18\%_{(0.09)}$ $0.00\%_{(0.00)}$ $26.92\%_{(0.08)}$ $5.13\%_{(0.01)}$ $0.00\%_{(0.00)}$
Eureka $90.00\%_{(0.00)}$ $83.89\%_{(0.28)}$ $90.56\%_{(0.39)}$ $\mathbf{92.78\%}_{(0.48)}$ $75.56\%_{(0.31)}$ $11.11\%_{(0.64)}$
AED $\mathbf{95.00\%}_{(0.04)}$ $\mathbf{87.77\%}_{(0.03)}$ $\mathbf{100\%}_{(0.00)}$ $\mathbf{90.00\%}_{(0.01)}$ $\mathbf{90.56\%}_{(0.10)}$ $\mathbf{12.22\%}_{(0.01)}$
Rear VDARS $2.28\%_{(0.03)}$ $1.49\%_{(0.01)}$ $1.94\%_{(0.01)}$ $1.45\%_{(0.01)}$ $0.85\%_{(0.01)}$ $0.00\%_{(0.00)}$
Eureka $53.33\%_{(0.72)}$ $58.33\%_{(0.36)}$ $13.89\%_{(0.15)}$ $90.56\%_{(0.44)}$ $47.22\%_{(0.39)}$ $36.11\%_{(0.28)}$
AED $\mathbf{60.56\%}_{(0.05)}$ $\mathbf{63.89\%}_{(0.06)}$ $\mathbf{31.11\%}_{(0.06)}$ $\mathbf{95.60\%}_{(0.01)}$ $\mathbf{82.22\%}_{(0.10)}$ $\mathbf{55.00\%}_{(0.23)}$
Roundabout Left VDARS $2.57\%_{(0.02)}$ $1.67\%_{(0.01)}$ $0.58\%_{(0.01)}$ $1.54\%_{(0.01)}$ $2.70\%_{(0.03)}$ $0.00\%_{(0.00)}$
Eureka $\mathbf{7.22\%}_{(0.08)}$ $5.00\%_{(0.08)}$ $8.89\%_{(0.20)}$ $1.11\%_{(0.15)}$ $1.11\%_{(0.07)}$ $0.00\%_{(0.00)}$
AED $\mathbf{6.67\%}_{(0.01)}$ $\mathbf{8.33\%}_{(0.03)}$ $\mathbf{8.89\%}_{(0.01)}$ $\mathbf{3.33\%}_{(0.01)}$ $\mathbf{4.44\%}_{(0.01)}$ $0.00\%_{(0.00)}$
Right VDARS $11.07\%_{(0.07)}$ $12.23\%_{(0.07)}$ $1.46\%_{(0.03)}$ $13.31\%_{(0.02)}$ $\mathbf{17.01\%}_{(0.19)}$ $0.00\%_{(0.00)}$
Eureka $\mathbf{12.78\%}_{(0.28)}$ $8.33\%_{(0.13)}$ $2.22\%_{(0.08)}$ $17.22\%_{(0.61)}$ $11.11\%_{(0.15)}$ $5.00\%_{(0.00)}$
AED $10.56\%_{(0.02)}$ $\mathbf{10.56\%}_{(0.01)}$ $\mathbf{3.89\%}_{(0.01)}$ $\mathbf{24.44\%}_{(0.04)}$ $\mathbf{16.11\%}_{(0.03)}$ $\mathbf{8.33\%}_{(0.01)}$
Rear VDARS $6.13\%_{(0.04)}$ $\mathbf{13.90\%}_{(0.08)}$ $5.26\%_{(0.09)}$ $10.74\%_{(0.01)}$ $7.29\%_{(0.08)}$ $0.00\%_{(0.00)}$
Eureka $3.33\%_{(0.36)}$ $2.78\%_{(0.08)}$ $15.56\%_{(0.43)}$ $5.56\%_{(0.20)}$ $7.78\%_{(0.20)}$ $\mathbf{5.56\%}_{(0.40)}$
AED $\mathbf{6.67\%}_{(0.01)}$ $3.33\%_{(0.03)}$ $\mathbf{21.67\%}_{(0.02)}$ $\mathbf{17.78\%}_{(0.05)}$ $\mathbf{11.67\%}_{(0.04)}$ $\mathbf{5.00\%}_{(0.01)}$
TABLE I: Effective Vulnerability Rates across different environments, tested policies, and numbers of adversarial vehicle agents. AED consistently achieves the highest effective vulnerability rates.
To evaluate the effectiveness of our method, we measure the effective vulnerability rate, which quantifies how frequently an adversarial policy induces a specific type of effective failure in the tested driving policy, where responsibility can be attributed to the tested policy. This metric represents the percentage of simulation runs where the adversarial policy successfully triggers such an effective failure. The results are shown in Table I.
AED achieves the highest effective vulnerability rates across a wide range of environments, tested policies, and numbers of adversarial vehicles, consistently outperforming or matching the baseline VDARS. It successfully discovers and exploits hidden weaknesses in the tested driving policies, including failure modes that VDARS fails to uncover. For instance, in the Highway environment with a VI-driven tested policy and two adversarial vehicle agents, AED achieves a $100\%$ effective failure rate for Left Lane Change, compared to only $34.24\%$ by VDARS. Similarly, in the Roundabout environment with a DQM-driven tested policy and two adversarial vehicle agents, AED improves the effective Rear-End failure rate from $5.26\%$ to $21.67\%$. These results demonstrate that AED reliably increases the effectiveness of induced failures across different traffic scenarios.
Ablation study: effect of preference-based enhancement
To assess the impact of the preference-based enhancement module, we compare AED with an ablated version, Parallel Eureka, which applies LLMs-driven reward functions without the final enhancement step. This isolates the impact of preference-based learning on reward quality and effectiveness. The results reveal that AED outperforms or matches Parallel Eureka in nearly all settings, and shows especially large gains when the original LLMs-driven rewards are only partially effective. For instance, in the Highway environment with a D3QN-driven tested policy and three adversarial vehicle agents, AED improves the effective Left Lane Change vulnerability rate improves significantly from $32.22\%$ with Parallel Eureka to $88.89\%$. This improvement highlights the impact of preference-based enhancement in guiding adversarial vehicles to consistently orchestrate failure scenarios that the tested policy struggles to handle. Additionally, in cases where Parallel Eureka already performed well, AED either maintains or improves performance, indicating that the preference-based refinement enhances reward fidelity without degrading strong initial signals.
These results confirm that AED is the most effective method for both uncovering hidden vulnerabilities in the tested driving policy and exploiting those weaknesses to trigger consistent failures. The preference-based enhancement plays a crucial role in converting suboptimal LLMs-driven reward signals into high-precision objectives for adversarial policy training.
Figure 4: Case study demonstrating the AED framework. (a) Task description for LLMs outlines the targeted Left Lane Change vulnerability, scenario details, and responsibility alignment. (b) The LLMs-driven reward function verifies the tested policy’s maneuver, ensures adversarial agents maintain non-aggressive behavior, and attributes fault appropriately. (c) Preference-based reward model differentiates the targeted and unintended scenarios.
V-D Interpretability and Discussion
To illustrate the functionality of the AED framework, we present a case study in Fig. 4, identifying Left Lane Change vulnerability in the Highway environment with a D3QN-driven tested policy and three adversarial vehicle agents. First, we prompt the LLMs with relevant environment context and task description, including the targeted accident type (Left Lane Change Collision), a textual description of traffic scenarios where the tested policy executes a left lane change maneuver that results in a collision, and responsibility criteria to determine the tested policy’s fault.
Based on these descriptions, the LLM-generated $K$ (In our experiment, $K$ was set to be $6$) executable reward functions for Left Lane Change Collision and trained $K$ corresponding agents in the environment using these reward functions. Next, we use these agents and the policy under evaluation to generate accident videos in the environment. Then, a human evaluator selects the best reward function based on the proportion of effective left-lane-change collisions caused by the agents in the videos.
The chosen reward function is presented in the middle of Fig. 4. We can observe that the generated reward function contains key components to ensure the function’s effectiveness. It first verifies that the tested AV’s maneuver prior to the collision was indeed a Left Lane Change. To prevent aggressive adversarial behavior, it also includes constraints that limit the adversarial vehicles’ steering angles and acceleration. Finally, the reward function determines if the collision can be attributed to the tested policy.
After adversarial training with the LLMs-driven reward function, the distribution of collected trajectories is as follows: Left Lane Change accidents $(36.67\%)$, Rear-End collisions $(1.00\%)$, and ineffective accidents $(62.33\%)$. Subsequently, training a reward model using preference-based learning enables a better differentiation between accident types, assigning higher rewards to Left Lane Change and lower rewards to unintended accidents. By incorporating both the preference reward model and the original LLMs-driven reward model into RL training, the proportion of effective Left Lane Change accidents significantly increases to $88.89\%$.
VI CONCLUSION
We propose AED, an LLM-based framework that uses reinforcement learning to automatically discover effective and diverse vulnerabilities of autonomous driving policies. We use an LLM to automatically design reward functions without the need of human experts. To enhance the diversity of vulnerabilities, we use the LLM the consider a broad range of potential accidents and train adversarial policies using RL in parallel. We further improve the effectiveness of each vulnerability by training a reward model with preference-based learning to filter ineffective accidents and other types of accidents. Experiments on various driving scenarios with both planning-based policies and learning-based policies show that AED discovers a diverse range of vulnerabilities and archives higher attack success rate compared to existing methods, which shows the potential of using LLMs to reduce the need for manual reward engineering and improve the comprehensiveness of autonomous driving safety evaluation.
ACKNOWLEDGMENT
The authors gratefully acknowledge the support from TOYOTA. This research was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496.